OpenStudy (shadow13):

i think it is c Which property is used in the statement? If y + 5 = 3, then y = –2. transitive property symmetric property addition property of equality subtraction property of equality

OpenStudy (calculusxy):

Why do you think that it is C?

OpenStudy (shadow13):

well i dont know i took a guess

OpenStudy (calculusxy):

So the addition property of equality would be used in a case like this: y - 5 = 3 because the inverse operation of subtraction would be addition.

OpenStudy (calculusxy):

Do you still think it's C?

OpenStudy (shadow13):

no now i think it is b is that right

OpenStudy (calculusxy):

I am not sure about the rest honestly. But here are some things that you should take into factor: Symmetric: If a = b, then b = a Transitive: If a = b and b = c, then a = c Subtraction property of equality: it would be used in a case like where you have to do the inverse operation of addition.

OpenStudy (calculusxy):

But may I know what grade you're in?

OpenStudy (calculusxy):

...so that I can help you accordingly?

OpenStudy (shadow13):

i am in 10th grade

OpenStudy (calculusxy):

Okay. So from the info that I gave you, what do you think you might want to cross out and why?

OpenStudy (shadow13):

i think i should cross out c and d because we already know that it is addition and there is no multiplication in the problem.

OpenStudy (calculusxy):

I have a feeling that it can be B the remaining answer. Explanation: a = b and b = a -2 + 5 = 3 , where (-2 + 5) is "a" and 3 is "b" -- one way 3 = -2 +5 -- second way

OpenStudy (shadow13):

ok well i know it is not a because b makes since on this question ill try that then

OpenStudy (anonymous):

https://www.youtube.com/watch?v=dQw4w9WgXcQ

OpenStudy (shadow13):

Which property is used in the statement? if k/2 =5 then k = 10 multiplication property of equality symmetric property reflexive property subtraction property of equality i thin k it is b but i change that i think it is c

OpenStudy (shadow13):

the reason that i think that is because its not symmetric and i don't see any multiplying or subtracting so thats y i think it is C

OpenStudy (calculusxy):

@shadow13 Sorry but I have to go now.

OpenStudy (shadow13):

ok

OpenStudy (shadow13):

@taylorclaudio

OpenStudy (shadow13):

@audiiii

OpenStudy (shadow13):

@JoKeR0331

OpenStudy (taylorclaudio):

lol thanksshadow

OpenStudy (shadow13):

welcome

OpenStudy (taylorclaudio):

b does make more cents looking at it now

OpenStudy (shadow13):

why is that